Overproduction of uric acid in the body can be an indication of cell breakdown. This may be an advance indication of illness such as gout, leukemia, or lymphoma.† Over a period of months, an adult male patient has taken thirteen blood tests for uric acid. The mean concentration was x = 5.35 mg/dl. The distribution of uric acid in healthy adult males can be assumed to be normal, with σ = 1.79 mg/dl. (a) Find a 95% confidence interval for the population mean concentration of uric acid in this patient's blood. What is the margin of error? (Round your answers to two decimal places.) lower limit upper limit margin of error (b) What conditions are necessary for your calculations? (Select all that apply.) σ is known n is large σ is unknown normal distribution of uric acid uniform distribution of uric acid (c) Interpret your results in the context of this problem. The probability that this interval contains the true average uric acid level for this patient is 0.95. We are 5% confident that the true uric acid level for this patient falls within this interval. We are 95% confident that the true uric acid level for this patient falls within this interval. The probability that this interval contains the true average uric acid level for this patient is 0.05. (d) Find the sample size necessary for a 95% confidence level with maximal margin of error E = 1.08 for the mean concentration of uric acid in this patient's blood. (Round your answer up to the nearest whole number.) blood tests
Answer #1

a. As population standard deviation is known we will use z distribution to find CI

z value for 95% CI is 1.96 as P(-1.96<z<1.96)=0.95

So Margin of Error is E=z*\frac{\sigma }{\sqrt{n}}=1.96*\frac{1.79}{\sqrt{13}}=0.97

Hence CI is CI=\overline{x} \pm E=5.35 \pm 0.97=(4.38,6.32)

Lower limit is 4.38

Upper limit is 6.32

Margin of Error is E=0.97

b. σ is known normal distribution of uric acid

c. We are 95% confident that the true uric acid level for this patient falls within this interval.

d. E=1.08

z value for 95% CI is 1.96 as P(-1.96<z<1.96)=0.95

So n=(\frac{z*\sigma}{E})^2=(\frac{1.96*1.79}{1.08})^2=10.55=11